# Identification of Semiparametric Panel Multinomial

### v2

@@ Appendix C Sufficient Conditions for Point Identification
Appendix D Additional Simulation Results
D.1 Adaptive-Grid Computation Algorithm
@@
In this section, we illustrate a typical output of our second-step computation algorithm based on the adaptive-grid search over the angle space, and show that the algorithm works well. For this purpose we consider a simplified DGP without fixed effect $A_{ij}$. We draw each of $X_{ijt,d}$ independently across each dimension $d$ from the standard normal distribution, and set the distribution of the idiosyncratic shock to be $\epsilon_{ijt}\sim TIEV\left(0,1\right)$, so that we can skip the first-step estimation and directly calculate the true conditional choice probability.. Note that the conditions for point identification of $\beta_{0}$ are satisfied. Because we are only seeking to illustrate the validity of the algorithm itself, we set $N$ to be large with $N=10^{7}$ and $D=3,\ J=3,\ T=2$. Then we apply our adaptive-grid algorithm to search for $\beta_{0}$.
@@
D.2 Estimation without Point Identification
[…]
@@ D.2 Estimation without Point Identification
(i) yes  - .0570 .0604 .0630 .0497 .0541 .0565
(ii) no  .01 .0585 .0605 .0640 .0505 .0538 .0566
 .1 .0514 .0732 .0841 .0438 .0692 .0780
 1 .0406 .1997 .2375 .0384 .1988 .2362
[…]
D.3 Results Varying $D,J,T$
@@
$A_{ij}\sim\begin{cases}0,&j=1,\\ \left[Z_{i}\right]_{+},&j=2,\\ \mathcal{U}\left[-0.25,0.25\right],&j=3,...,J,\end{cases}$ $X_{ijt,d}\sim\begin{cases}U\left[-1,1\right],&d=1,\\ Z_{i}+\mathcal{N}\left(0,6\right),&d=2,\\ \mathcal{N}\left(0,1\right),&d=3,...,D,\end{cases}$
which coincides with the baseline model at $D=3,\ J=3$. We emphasize that in all configurations we allow for nonlinear dependence between $A$ and $X$ via the latent variable $Z_{i}$.
@@
$D=3$0522 .0357 .1142 .0877
$D=4$0650 .0467 .1325 .0984
From Table 8 we find a larger $T$ improves the performance of our estimator, which is arguably more practically relevant given the increasing availability of long panel data. The improvement in performance with larger $T$ is because our method can extract more information from $T\times\left(T-1\right)$ ordered pairs of time periods which effectively increase the total number of observations. We also find that increase in $D$ or $J$ adversely affects the performance of our estimator, which is expected because more information is required to estimate more covariates or deal with more alternatives. For example, when $J$ is 3 and $T$ is 4, an increase in the dimension of product characteristics $D$ from 3 to 4 increases the rMSE from .0407 to .0488. The change in performance for increasing $J$ is more significant. For instance, when $D=4$ and $T=4$, an increase in $J$ from 3 to 4 increases the MND from .0467 to .0984.

### v3

@@ Appendix C Sufficient Conditions for Point Identification
Appendix D More on the Adaptive-Grid Algorithm
We elaborate on the adaptive-grid algorithm described in Section 3.3 and provide additional details regarding its practical implementation. A graphical illustration of the algorithm is presented in Figure 1.
  1. 1.
    Initialization and Coarse Grid Search. The search begins by defining a coarse, uniform grid of $M_{0}^{D-1}$ points over the initial parameter space $\Theta$. The code sets $M_{0}$ via the M_Step parameter. The objective function $Q(\beta)$ is evaluated at each point on this grid. This stage then iteratively refines the search space. After each evaluation, the algorithm prunes the parameter space by retaining only the region containing the points that yielded the lowest 20% of objective function values. A new, equally-sized grid is then constructed within this smaller, more promising region, and the process repeats. This iterative pruning continues for a fixed number of loops or until the minimum objective function value stabilizes, effectively narrowing the search to a region likely to contain the global minimum. The procedure includes a check to recenter the search space if the minimum appears to lie on a boundary of the angular coordinates, which prevents erroneously discarding a wrapped-around solution.
  2. 2.
    Fine Grid Refinement. Upon completion of the coarse search, the algorithm enters a refinement stage. It constructs a new, finer grid within a small neighborhood surrounding the set of candidate points that produced the minimum objective function value in the previous stage. The resolution of this grid is determined by a smaller step size, Tol_Step. By re-evaluating the objective function on this denser set of points, this stage seeks to improve upon the initial estimate. If a new, lower minimum value is found, the process repeats around this new set of points. This stage concludes when no further improvement in the minimum of the objective function can be achieved, indicating that the algorithm has converged to a local minimum within the refined region.
  3. 3.
    Boundary Identification. The final stage aims to precisely delineate the identified set of minimizing parameters. A very fine grid is constructed in the immediate buffer zone around the set of optimal points found in the prior stage, while excluding the interior of this set. This allows the algorithm to meticulously probe the boundary of the solution set. Any points on this boundary that also yield the minimum objective function value are added to the solution set. This step repeats, further reducing the grid step size until a pre-specified tolerance (Tol) is met and a clear separation exists between the parameters that minimize the function and those that do not. This ensures the final identified set is not only optimal but also well-defined.
Figure 1: An Adaptive-Grid Algorithm
[…]
Figure 1 provides a graphical illustration of our adaptive-grid algorithm to find the minimizer of the criterion function over the parameter space $\Theta$. The figure corresponds to a re-parameterization in the 2-dimensional angle space of a 3-dimensional unit ball, as specified in equation (16) of the paper for $D=3$. The pink area represents the initial angle space. The light blue rectangle corresponds to $\underline{\Theta}^{(1)}$ defined in (18). The dark blue rectangle represents the identified set $\Theta_{0}$, which is the best we can hope for. The disconnectedness of the light blue rectangle corresponds to Step 3 of the algorithm presented in Section 3.3. This is largely a programming nuisance: it means that if $\underline{\Theta}^{(1)}\not\subset\Theta^{(1)}$ due to the special structure of the angle space, we add $2\pi$ to $\underline{\Theta}^{(1)}$ such that it is connected and lies within $\Theta^{(1)}$. Note that adding $2\pi$ to the first coordinate of $\theta$ does not change the value of $\theta$ nor $\beta$.
Appendix E Counterfactual Analysis in Long Panels
So far, we have focused on the identification and estimation of the index parameter $\beta_{0}$. While $\beta_{0}$ itself contains rich information about the unobserved preference of consumers, we are often also interested in counterfactual parameters defined as some functional of not only $\beta_{0}$, but also other unknown components of the model. In this section, we discuss how the estimate $\hat{\beta}$ of $\beta_{0}$, and the computed indices based on $\hat{\beta}$, can be used to estimate more sophisticated counterfactual parameters.
An important class of counterfactual parameters concerns the prediction of counterfactual market shares (aggregate choice probability) in the form of
$\mu_{j}\left(\overline{{\bf X}}\right):=\int\mathbb{E}\left[\left.y_{ijt}\right|{\bf X}_{it}=\overline{{\bf X}},{\bf A}_{i}\right]\text{d}\mathbb{P}\left({\bf A}_{i}\right),$ (43)
Demand elasticities may be further computed as $\nabla\mu_{j}\left(\overline{{\bf X}}\right)$, which gives the marginal effect of an exogenous change in certain observable characteristics on consumer choices. [^24] It is important to note that in the expression of $\mu$ we use the marginal distribution $\mathbb{P}\left({\bf A}_{i}\right)$ rather than the conditional distribution $\mathbb{P}\left(\left.{\bf A}_{i}\right|{\bf X}_{it}=\overline{{\bf X}}\right)$. This separation between the exogenously imposed counterfactual $\overline{{\bf X}}$ and the distribution of the unobserved ${\bf A}_{i}$ is key to the interpretation of $\mu_{j}\left(\overline{{\bf X}}\right)$ as the direct effect of the exogenous change in observable characteristics $\bf X$} on choice probabilities, with the unobserved heterogeneity $\bf A$} unaffected by this exogenous change held fixed.
To achieve this separation, we seek to identify and estimate the integrand $\mathbb{E}\left[\left.y_{ijt}\right|{\bf X}_{it}=\overline{{\bf X}},{\bf A}_{i}\right]$, which is a function of $\beta_{0}$. Conditional on $\bm{\delta}_{it}=\overline{{\bf X}}^{{}^{\prime}}\beta_{0}=:\overline{\bm{\delta}}$ and ${\bf A}_{i}$, by our model specification (1) we have
$\mathbb{E}\left[\left.y_{ijt}\right|\bm{\delta}_{it}=\overline{\bm{\delta}},{\bf A}_{i}\right]=\psi_{j}\left(\overline{\bm{\delta}},\ {\bf A}_{i}\right)=:\psi_{ij}\left(\overline{\bm{\delta}}\right).$
Here, an important observation is that although the individual heterogeneity ${\bf A}_{i}$ is not directly observable, the identity of $i$ is observable. We can hold individual $i$ fixed in the regression to control for ${\bf A}_{i}$ and only use variations in the data across $t$ in a long panel setting to estimate the conditional choice probability. Specifically, suppose we have long panels, i.e., $T\rightarrow\infty$. Suppose the conditions in Appendix C are also satisfied so that $\delta_{0}$ is point identified. We can identify $\psi_{ij}\text{$\left(\overline{\bm{\delta}}\right)$}$ for each fixed $i$ and product $j$ and estimate it via nonparametric regression of $y_{ijt}$ on $\left(\delta_{i1t},\ldots,\delta_{iJt}\right)$ across $t=1,\ldots,T$, under suitable stationarity and weak dependence conditions on the error terms. See, for example, Su and Jin (2012) for a sieve estimator of functional fixed effects in a long panel setting.
[…]
Appendix F Additional Simulation Results
F.1 Adaptive-Grid Computation Algorithm
@@
In this section, we illustrate a typical output of our second-step computation algorithm based on the adaptive-grid search over the angle space, and show that the algorithm works well. For this purpose we consider a simplified DGP without fixed effect $A_{ij}$. We draw each of $X_{ijt,d}$ independently across each dimension $d$ from the standard normal distribution, and set the distribution of the idiosyncratic shock to be $\epsilon_{ijt}\sim TIEV\left(0,1\right)$, so that we can skip the first-step estimation and directly compute the true conditional choice probability. Note that the conditions for point identification of $\beta_{0}$ are satisfied. Because we are only seeking to illustrate the validity of the algorithm itself, we set $N$ to be large with $N=10^{7}$ and fix D=3,J=3,and T=2D=3,\ J=3,\ $\text${and }T=2. Then, we apply our adaptive-grid algorithm to search for $\beta_{0}$.
@@
F.2 Estimation without Point Identification
We now investigate the performance of our estimator when point identification fails. To make things comparable, we fix $\left(N,D,J,T\right)$ as in the baseline case, but modify the configuration in two different ways. We maintain the point identification in one setting but lose the point identification in the other. Specifically, we set $Z_{i}\sim\mathcal{U}\left[-\sqrt{3},\sqrt{3}\right]$, $X_{ijt,1}\sim\mathcal{U}\left[-1,1\right]$, $X_{ijt,2}=Z_{i}+\mathcal{N}\left(0,6\right)$, and $X_{ijt,3}\sim\mathcal{N}\left(0,1\right)$ for the point identified case. For the DGP without point identification, we let $Z_{i}\sim\mathcal{U}\left[-\sqrt{3},\sqrt{3}\right]$, $P(X_{ijt,1}=1/2)=P(X_{ijt,1}=-1/2)=0.5$, $X_{ijt,2}=Z_{i}+\mathcal{U}\left[-\sqrt{6},\sqrt{6}\right]$, and $P(X_{ijt,3}=1/2)=P(X_{ijt,3}=-1/2)=0.5$. The construction of $A_{ij}$ is the same as in the DGP for the baseline results in Table 1. We deliberately control the location and scale of each variable to be comparable across the two configurations, with the only differences being the presence of discreteness and boundedness of supports. When point identification fails, we compute the set estimator $\hat{B}_{\hat{c}}$ of (16) with $\hat{c}>0$. Table 6 contains simulation results under the two configurations, with different choices of $\hat{c}$ when point identification fails. [^25] Specifically, noting that $c_{N}\log N\leq N^{-1/4}\log N\approx 0.92\leq 1$ for $N=10,000$, we set $\hat{c}=0.01,$ $0.1$ and $1$, respectively.
@@ F.2 Estimation without Point Identification
(i) yes  - 0.0574 0.0636 0.0665 0.0501 0.0577 0.0602
(ii) no  0.01 0.0624 0.0657 0.0711 0.0536 0.0590 0.0630
 0.1 0.0521 0.0820 0.0966 0.0444 0.0782 0.0904
 1 0.0476 0.2198 0.2631 0.0458 0.2190 0.2619
Across rows in (i) and (ii), we see that the lack of point identification does negatively affect the performance of our estimates, but the impact is limited to a moderate degree. Within rows in (ii), we observe that, as expected, a more conservative choice of the constant $\hat{c}$ worsens performance of the upper and lower bounds by enlarging the estimated sets; meanwhile, it appears that the size and the performance of our estimator based on $\hat{\beta}^{m}$ is not sensitive to the choice of $\hat{c}$.
F.3 Results Varying $D,J,T$
@@
$A_{ij}\sim\begin{cases}0,&j=1,\\ \left[Z_{i}\right]_{+},&j=2,\\ \mathcal{U}\left[-0.25,0.25\right],&j=3,\ldots,J,\end{cases}$ $X_{ijt,d}\sim\begin{cases}U\left[-1,1\right],&d=1,\\ Z_{i}+\mathcal{N}\left(0,6\right),&d=2,\\ \mathcal{N}\left(0,1\right),&d=3,\ldots,D,\end{cases}$
which coincides with the baseline model at D=3,J=3D=3,\ J=3. We emphasize that in all configurations we allow for nonlinear dependence between $A$ and $X$ via $Z$.
@@
$D=3$ 0.0522 0.0363 0.1154 0.0860
$D=4$ 0.0660 0.0475 0.1264 0.1003
From Table 7 we find a larger $T$ improves the performance of our estimator, which is arguably more practically relevant given the increasing availability of long panel data. The improvement in performance with larger $T$ is because our method can extract more information from $T\times\left(T-1\right)$ ordered pairs of time periods which effectively increase the total number of observations. We also find that increase in $D$ or $J$ adversely affects the performance of our estimator, which is expected because more information is required to estimate more covariates or deal with more alternatives. For example, when $J$ is 3 and $T$ is 4, an increase in the dimension of product characteristics $D$ from 3 to 4 increases the rMSE from 0.0411 to 0.0497. The change in performance for increasing $J$ is more significant. For instance, when $D=4$ and $T=4$, an increase in $J$ from 3 to 4 increases the MND from 0.0475 to 0.1003.
F.4 Robustness Against the Blue-Bus/Red-Bus Problem
We now demonstrate the robustness of our method against the Blue-Bus/Red-Bus problem, or more precisely, a case where the product set includes highly substitutable alternatives.
Specifically, we modify the DGP in our baseline simulation by increasing $J$ from 3 to 4, and introduce the fourth product as an indistinguishable duplicate of the third one: they share the same observable and unobservable characteristics, including the observable characteristics $X$, unobserved fixed effect $A$, and the unobserved error term $\epsilon$. We equally split the market share between $j=3$ and $j=4$.
Using the above DGP, we compare our method’s performance to that of the RCLM estimator. For the RCLM estimator, we simulate $M=1,000$ random coefficients $\beta_{i}=\beta+\eta_{i}$ where $\eta_{i}\sim N\left(0,I_{D}\right)$ for the consumers with heterogeneous preferences, where $I_{D}$ is a $D\times D$ identity matrix. The predicted market shares are thus computed as
$\sigma_{jt}\left(X_{i\cdot t};\beta\right)$ $=\mathbb{E}\left[\left.\mathbf{\mathbbm{1}}\left\{U_{ijt}>\max_{k\neq j}U_{ikt}\right\}\right|X_{i\cdot t}\right]$
$=\int\dfrac{\exp\left(X_{ijt}^{{}^{\prime}}\beta+X_{ijt}^{{}^{\prime}}\eta_{i}\right)}{\sum_{k=1}^{J}\exp\left(X_{ikt}^{{}^{\prime}}\beta+X_{ikt}^{{}^{\prime}}\eta_{i}\right)}dF_{\eta}\left(\eta_{i}\right)$
$\simeq\dfrac{1}{M}\sum_{m=1}^{M}\dfrac{\exp\left(X_{ijt}^{{}^{\prime}}\beta+X_{ijt}^{{}^{\prime}}\eta_{i}^{\left(m\right)}\right)}{\sum_{k=1}^{J}\exp\left(X_{ikt}^{{}^{\prime}}\beta+X_{ikt}^{{}^{\prime}}\eta_{i}^{\left(m\right)}\right)}.$ (44)
For each round of simulation, we compute predicted market shares by numerically integrating over the random taste distribution over $M$ agents. We then choose the candidate $\beta$ over a fine grid of $L=10,000$ points from $\mathbb{\mathbb{S}}^{D-1}$ that achieves the smallest distance in the Euclidean norm between the predicted and observed market shares as $\hat{\beta}$. We use $N=1,000$, $D=3$, $J=4$, $T=2$, and B=B=1,000 for this simulation exercise.
Table 8 summarizes the results. The sum of absolute biases of $\hat{\beta}$ under our method is 0.0452, which is markedly lower than that of the RCLM estimator (0.8689). The average size of the estimated sets is 0.0585 for our method. Since, by construction, the RCLM estimator is a point estimator, its average size is zero across simulations. Finally, the rMSE and MND are significantly lower for our estimator than for the RCLM. The results show that, in the presence of highly similar products (with highly correlated unobserved heterogeneity), the RCLM estimator cannot recover the structural parameters accurately. In contrast, our method is robust to arbitrary dependence structures in the unobserved heterogeneity terms across products, leading to more accurate estimates of the structural parameters.
Table 8: Illustration of the “Blue-Bus/Red-Bus” Problem
$\sum_{d}\left|\text{bias}_{d}\right|$ $\sum_{d}\text{mean(u-l)}_{d}$ rMSE MND
Current Paper 0.0452 0.0585 0.1389 0.1210
RCLM 0.8689 - 0.6047 0.6042